2gram of o2 at 27 c0 and 780 mm of Hg pressure volume ?

Grade:12th pass

1 Answers

Arun
6 years ago
Formula used to calculate volume :
PV = nRT
​V = n RT /P.....(1)
n = number of moles of Oxygen
  = given mass of oxygen/molar mass of oxygen
 = 2 g /32 g mol-1 = 0.0625
​R = 0.0821 L atm K-1 mol-1
​T = 27° C = (27 + 273) K = 300 K
​P = 780 mm Hg = 1.03 atm     .....(760 mm Hg = 1atm)
Put the given values in eq. (1)
 
You will get Volume.
 
 
Regards
Arun (askIITians forum expert)